21 In binomial distribution ‘n’ number of trials are
22 In binomial distribution range of values of variable x is
23 In normal distribution the mean, median, mode
24In which of the following binomial distributions is the normal approximation appropriate
25 In which probability distribution mean and variance are equal
26 Maximum ordinate of normal distribution when µ = 50, σ= 6, and x = 50 is equal to
27 Mean of binomial distribution is
1
Expert's answer
2020-07-06T20:36:29-0400

21)There are a fixed number of trials, n. There are only two possible outcomes, called “success” and, “failure” for each trial. The letter p denotes the probability of a success on one trial and q denotes the probability of a failure on one trial.The n trials are independent and are repeated using identical conditions.

22)In binomial distribution range of values of variable x is the range of values that the random variable X may assume, might be from 0 to n

23)The mean, median, and mode of a normal distribution are equal.

24) The normal distribution can be used as an approximation to the binomial distribution, under certain circumstances, namely: If X ~ B(n, p) and if n is large and/or p is close to ½, then X is approximately N(np, npq)

25) For the Poisson distribution the mean and variance are always equal.

26)Max(f(x)):d/dx(f(x))=-2(x-µ)f(x)

for d/dx(f(x))=0

x=µ

At which f(x)=1/σ(2pi)0.5

27) The mean of the distribution (μx) is equal to n * P